the diagonals of rhombus FGHJ intersect at point K. If side GH is equal to 3x-10 side JH is equal to 6x-19, find x.
Answer by nyc_function(2741)

Rhombus
I have all of the properties of the parallelogram PLUS
- 4 congruent sides
- diagonals bisect angles
- diagonals perpendicular

3x - 10 = 6x - 19

3x - 6x = -19 + 10

-3x = -9

x = -9/-3

x = 3